# Scaling Laws of RoPE-based Extrapolation

### v1

@@ A.1 RoPE from Sequence Domain
While RoPE can theoretically convey the relative information at any context length, RoPE still fails to extrapolate practically. It is worth noting that rotary angles $\bm{\theta}$ in Equation 9 play an important role. In the vanilla design of RoPE, $\bm{\theta}$ is defined as Equation 10. Different angles correspond to different features and that is the starting point of most RoPE-based extrapolation methods (LocalLLaMA, 2023b; Rozière et al., 2023; Pal et al., 2023) shown in Table 2.
@@ A.2 RoPE from Frequency Domain
Name Applying Phase Training Length Context Window Ratio Auxiliary Window Note
xPos (Sun et al., 2022) Pre-training 1K 4K ×$\times 4$ BCA
Linear PI (Chen et al., 2023) Fine-tuning 16K 16K ×$\times 1$
Giraffe (Pal et al., 2023) Fine-tuning 4K 32K ×$\times 8$ truncated
Code LLaMA (Rozière et al., 2023) Fine-tuning 16K 100K+ ×$\times 6$
YaRN (Peng et al., 2023) Fine-tuning 64K 128K ×$\times 2$ window ppl
Fixed NTK (LocalLLaMA, 2023b) Inference 4K 16K ×$\times 4$
Dynamic NTK (LocalLLaMA, 2023a) Inference 4K 16K+ ×$\times 4$
PCW (Ratner et al., 2022) Inference 2K 6K ×$\times 3$ for ICL
NBCE (Su, 2023a) Inference 2K 10K ×$\times 5$ for ICL
LM-Infinite (Han et al., 2023) Inference 4K 32K+ ×$\times 8$ $\Lambda$-shaped
Scaling RoPE (Ours) Fine-tuning 4K 16K 100K+ 1M+ ×$\times 25$$\bm{\times 64}$ base=1000000 base=500
@@ B.1 Experiment Setup
We conduct experiments on the pre-trained 7B and 13B LLaMA2 models (Touvron et al., 2023b). For fine-tuning 7B and 13B models, we use 32 A100 GPUs and adopt ZeRO3 strategies (Rajbhandari et al., 2020). We use AdamW (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We set the learning rate to $2\times 10^{-5}$ with no warmup. We set the max gradient norm to 2.5 for 7B and 1 for 13B respectively. We set the weight decay to zero.
For fine-tuning RoPE with different bases, we set the global batch size to 128, tuning the context length to 4K, the same as the training length, and the evaluating context length to 100K. We fine-tune the models for 1K steps using the next token prediction objective with training data from the Pile (Gao et al., 2020) and compare the tuning performance on the validation set of Books3 subset (Presser, 2020) from the Pile. We fine-tune LLaMA2 with CoLLiE [^3] https://github.com/OpenLMLab/collie, a collaborative toolbox for tuning large language models in an efficient way, and conduct evaluation discussed in Appendix B.2 with OpenCompass [^4] https://opencompass.org.cn/. Both training and testing are accelerated by FlashAttention-2 (Dao, 2023).
We compare results with mainstream extrapolating strategies, such as Linear Position Interpolation (Chen et al., 2023) and NTK method (LocalLLaMA, 2023b; a). For fine-tuning with Linear PI, we set the global batch size to 64 and tuning length to 8K, which follows Chen et al. (2023). For fine-tuning with 16K context length, wet set the global batch size 32. When we fine-tune LLaMA2 7B with the last 36 dimensions being cut off, discussed in Section 3.4, we set the softmax scale $\sqrt{92}$, the square root of updated dimension size of $\bm{q}_{t},\bm{k}_{s}$, and keep other setups the same. Except for the position embedding, we do not modify the architecture of LLaMA2 (Touvron et al., 2023b).
@@ B.2 Short Context Validation
Next, we validate whether Scaling RoPE, namely the fine-tuning RoPE with different bases on original context length, has a side effect on LLM. We use short context tasks, such as NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), MMLU (Hendrycks et al., 2020) and SuperGLUE (Wang et al., 2019), to assess whether the model loses the knowledge acquired in pre-training after the fine-tuning process. The results obtained are shown in the Table 3. We find that the fine-tuning results, though slightly lower compared to the original LLaMA2 7B and 13B (Touvron et al., 2023b) and directly fine-tuned version, are still acceptable. In fact, in some tasks, the performance of Scaling-RoPE even appears to be better than before. This indicates that our approach not only enhances the extrapolation performance but also does not significantly harm the inherent knowledge of LLM.
NQ TriviaQA MMLU SuperGLUE
LLaMA2 7B (Touvron et al., 2023b) 19.06 54.09 46.78 52.99
NTK fixed $\alpha=8$ (LocalLLaMA, 2023b) 19.89 53.16 39.36 51.84
Fine-tuned 20.14 55.14 47.73 55.68
Linear PI $\lambda=2$ (Chen et al., 2023) 18.59 53.32 47.27 54.71
Scaling RoPE base=2608 16.01 51.21 37.07 53.75
Scaling RoPE base=1304 17.70 51.27 33.46 54.55
Scaling RoPE base=652 17.12 48.90 33.43 54.47
Scaling RoPE base=500 20.36 52.43 38.70 53.42
Scaling RoPE base=40000 17.34 55.27 47.34 54.85
Scaling RoPE base=80000 17.81 54.72 46.87 54.88
Scaling RoPE base=160000 18.98 55.28 46.24 55.02
Scaling RoPE base=400000 20.39 54.91 46.01 55.55
Scaling RoPE base=600000 19.20 54.75 46.39 54.77
Scaling RoPE base=1000000 19.50 54.76 46.29 55.09
LLaMA2 13B (Touvron et al., 2023b) 24.82 60.69 55.77 54.22
NTK fixed $\alpha=8$ (LocalLLaMA, 2023b) 22.77 60.81 52.04 53.99
Fine-tuned 26.68 61.57 55.78 55.86
Linear PI $\lambda=2$ (Chen et al., 2023) 25.96 60.78 54.37 53.58
Scaling RoPE base=2608 20.00 59.62 55.06 55.93
Scaling RoPE base=1304 13.60 59.12 53.59 56.14
Scaling RoPE base=652 15.98 58.67 50.49 56.54
Scaling RoPE base=500 11.50 57.54 47.98 56.71
Scaling RoPE base=40000 26.87 60.80 55.26 55.77
Scaling RoPE base=80000 26.76 61.25 55.32 55.55
Scaling RoPE base=160000 27.06 61.22 54.35 55.41
Scaling RoPE base=400000 27.53 60.94 54.01 55.04
Scaling RoPE base=600000 26.81 61.28 53.75 55.69
Scaling RoPE base=1000000 26.51 59.22 52.94 55.33
@@
Table 3: Short context validation results of scaling RoPE with different base and other extrapolating methods in LLaMA2 (Touvron et al., 2023b). The first box is the result of 7B size and the second box is those of 13B version. LLaMA2 7B or 13B corresponds to the original result as well as that of Dynamic NTK (LocalLLaMA, 2023a) since Dynamic NTK does not affect the model within the training context. For fixed NTK (LocalLLaMA, 2023b), $\alpha=8$, while for linear position interpolation, $\lambda=2$, the ratio between the fine-tuning context length and the pre-training context length. The final score of MMLU and SuperGLUE is the average score of all corresponding subtasks.
@@ Appendix D Discussion
Besides, we discuss the instructive value of our theory for other extrapolation strategies focused on achieving longer context during the testing phase. These methods are still necessary given two facts. On one hand, the performance of RoPE with a smaller base is still left behind compared with RoPE with much larger bases, such as 1000000, as shown in Figure 1. On the other hand, for RoPE with a base that is not large enough, it still can not extrapolate to a context of 100K or longer as shown in Figure 2. In order to further enhance RoPE’s adaptability to a longer context, whatever the base value is, we discuss the effect of two complementary methods in the inference phase, log-scaled attention (Su, 2023b) and dynamic-scaled RoPE (LocalLLaMA, 2023a) on RoPE with different bases.
@@
Besides Log-scaled attention, window method, such as sliding window and its variant, is also a widely accepted strategy for extrapolation, used in inference or evaluation (Press et al., 2021; Sun et al., 2022). Compared with the above strict window-based method, we follow the xPos method proposed in Sun et al. (2022), shown in Equation 18, originally used in the pre-training phase. In this work, we regard this method as a soft sliding window used in the inference phase as use it as a further complement to the log-scaled method. Still, we do little modification besides using the $T_{\text{extra}}$ as the denominator instead of the original denominator $T_{\text{train}}$.
@@
namely Dynamic NTK (LocalLLaMA, 2023a) is a widely used extrapolation. Here, we only do two little modifications. One is to change the base 10000 in vanilla RoPE (Su et al., 2021) with the base scaled in the fine-tuning phase, $\beta$. The other is still to replace the $T_{\text{train}}$ in Equation 7 with $T_{\text{extra}}$ we derive given Equation 14 in Theorem 3.
for $\text{for }\bm{A}_{t,s}\text{, }\theta_{n}={\left(\beta\cdot\alpha_{t}\right)}^{-2n/d}\text{, where }\alpha_{t}=\max\left(1,2^{\left\lceil\log_{2}{\frac{t}{T_{\text{extra}}}}\right\rceil+1}-1\right)\text{.}$ (19)
@@
Figure 11: Perplexity on the validation data from Books3 dataset (Presser, 2020) of LLaMA 7B (Touvron et al., 2023b) based on RoPE with different bases enhanced with log-scaled attention, xPos and dynamic-scaled RoPE and their combination. Here, wrong means calculating $\alpha_{t}$ in Dynamic NTK (LocalLLaMA, 2023a) or $p_{t}$ in (Su, 2023b) with training length $T_{\text{train}}$, while correct means using extrapolation upper bound $T_{\text{extra}}$ instead of $T_{\text{train}}$.

### v2

@@ 5 Conclusion
Acknowledgements
This work is supported by Shanghai Artificial Intelligence Laboratory. Special thanks to Qipeng Guo, Shuo Zhang, and Kai Lv, whose invaluable suggestions have greatly contributed to this work.
@@ A.1 RoPE from Sequence Domain
While RoPE can theoretically convey the relative information at any context length, RoPE still fails to extrapolate practically. It is worth noting that rotary angles $\bm{\theta}$ in Equation 8 play an important role. In the vanilla design of RoPE, $\bm{\theta}$ is defined as Equation 9. Different angles correspond to different features and that is the starting point of most RoPE-based extrapolation methods (bloc97, 2023b; Rozière et al., 2023; Pal et al., 2023) shown in Table 2.
@@
For example, the NTK method (bloc97, 2023b; a) reduces the rotary angle under the maintenance of its value form, thus enabling RoPE-based LLM to adapt to position information in a longer context without further training.
@@ A.2 RoPE from Frequency Domain
Name
Applying
Phase
Training
Length
Context
Window
Ratio
Auxiliary
Window
Note
xPos
(Sun et al., 2023)
Pre-training 1K 4K ×4$\times$ 4× 4 BCA
Linear PI
(Chen et al., 2023)
Fine-tuning 16K 16K ×1$\times$ 1× 1
Giraffe
(Pal et al., 2023)
Fine-tuning 4K 32K ×8$\times$ 8× 8 truncated
Code LLaMA
(Rozière et al., 2023)
Fine-tuning 16K 100K+ ×6$\times$ 6× 6
YaRN
(Peng et al., 2023)
Fine-tuning 64K 128K ×2$\times$ 2× 2 window ppl
Fixed NTK
(bloc97, 2023b)
Inference 4K 16K ×4$\times$ 4× 4
Dynamic NTK
(bloc97, 2023a)
Inference 4K 16K+ ×4$\times$ 4× 4
PCW
(Ratner et al., 2022)
Inference 2K 6K ×3$\times$ 3× 3 for ICL
NBCE
(Su, 2023a)
Inference 2K 10K ×5$\times$ 5× 5 for ICL
LM-Infinite
(Han et al., 2023)
Inference 4K 32K+ ×8$\times$ 8× 8 $\Lambda$-shaped
Scaling RoPE
Fine-tuning
4K
16K
100K+
1M+
×25$\times$ 25× 25
×$\bm{\times 64}$
base=1000000
base=500
@@ B.1 Experiment Setup
We conduct experiments on the pre-trained 7B and 13B LLaMA2 models (Touvron et al., 2023b). For fine-tuning 7B and 13B models, we use 32 A100 GPUs and adopt ZeRO3 strategies (Rajbhandari et al., 2020). We use AdamW (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We set the learning rate to $2\times 10^{-5}$ with no warmup. We set the max gradient norm to 2.5 for 7B and 1 for 13B respectively. We set the weight decay to zero.
For fine-tuning RoPE with different bases, we set the global batch size to 128, tuning the context length to 4K, the same as the training length, and the evaluating context length to 100K. We fine-tune the models for 1K steps using the next token prediction objective with training data from the Pile (Gao et al., 2021) and compare the tuning performance on the validation set of Books3 subset (Presser, 2020) from the Pile. We fine-tune LLaMA2 with CoLLiE(Lv et al., 2023), a collaborative toolbox for tuning large language models in an efficient way, and conduct evaluation discussed below with OpenCompass [^3] https://opencompass.org.cn/. Both training and testing are accelerated by FlashAttention-2 (Dao, 2023).
We compare results with mainstream extrapolating strategies, such as Linear Position Interpolation (Chen et al., 2023) and NTK method (bloc97, 2023b; a). For fine-tuning with Linear PI, we set the global batch size to 64 and tuning length to 8K, which follows Chen et al. (2023). For fine-tuning with 16K context length, wet set the global batch size 32. When we fine-tune LLaMA2 7B with the last 36 dimensions being cut off, discussed in Section 3.4, we set the softmax scale $\sqrt{92}$, the square root of updated dimension size of $\bm{q}_{t},\bm{k}_{s}$, and keep other setups the same. Except for the position embedding, we do not modify the architecture of LLaMA2 (Touvron et al., 2023b).
@@ B.2 Short Context Validation
Lambada SuperGLUE Hellaswag PIQA Winogrande OBQA
LLaMA2 7B 73.30 50.43 69.45 76.66 61.25 32.00
NTK fixed $\alpha=8$ 67.53 51.13 68.43 76.28 61.48 29.40
Linear PI $\lambda=2$ 72.64 50.71 69.75 76.01 61.56 40.60
Linear PI $\lambda=4$ 71.82 52.34 69.48 76.33 61.56 30.40
base=500 69.51 48.34 66.46 74.97 59.67 31.00
base=652 68.78 46.86 64.00 75.35 60.22 28.80
base=1304 69.24 47.22 65.51 76.06 60.30 25.00
base=2608 69.38 48.04 65.61 75.24 60.93 28.00
base=10000 72.62 51.49 69.60 75.95 61.40 38.00
base=40000 72.77 51.02 69.46 76.06 61.33 36.20
base=80000 72.68 51.20 69.44 76.01 60.69 37.40
base=160000 72.54 50.24 69.07 76.01 60.77 36.80
base=400000 72.25 51.08 68.86 75.68 60.85 36.00
base=600000 72.09 51.02 68.90 75.63 61.33 34.80
base=1000000 71.80 50.84 68.82 75.63 60.77 35.40
base=500, 16K 68.62 48.20 62.35 73.67 58.33 25.00
base=10000, 16K 72.17 50.60 67.68 75.79 60.22 28.80
base=40000, 16K 70.77 48.03 67.69 76.33 61.33 28.20
base=80000, 16K 72.48 51.29 69.54 76.33 61.09 32.60
base=120000, 16K 72.66 50.34 69.54 76.22 61.96 32.20
base=1000000, 16K 71.14 49.80 68.45 75.90 61.88 29.80
LLaMA2 13B 76.48 52.35 74.22 79.54 63.85 42.60
NTK fixed $\alpha=8$ 72.56 52.64 73.56 79.71 62.67 37.00
Linear PI $\lambda=2$ 75.18 63.15 74.93 78.94 63.69 51.40
Linear PI $\lambda=4$ 74.29 61.64 74.81 79.38 63.77 50.60
base=500 72.56 52.81 71.38 76.88 64.17 35.20
base=652 73.78 52.54 71.62 77.48 62.83 35.20
base=1304 74.21 52.19 72.38 77.37 64.01 36.80
base=2608 74.54 52.54 73.63 78.02 63.22 43.40
base=10000 75.12 53.19 75.19 79.00 62.98 55.00
base=40000 74.87 53.11 74.76 79.60 62.98 52.60
base=80000 74.89 52.37 74.76 79.27 62.90 53.80
base=160000 74.89 52.63 74.50 79.05 62.75 54.20
base=400000 74.23 52.47 74.50 78.89 62.19 54.40
base=600000 74.19 54.14 74.24 79.00 63.77 53.60
base=1000000 74.58 50.82 74.36 78.67 62.98 51.80
base=500, 16K 69.34 49.28 67.23 76.50 58.48 29.60
base=10000, 16K 74.00 53.44 73.16 78.89 63.93 35.00
base=40000, 16K 75.06 54.08 74.33 78.78 62.51 38.00
base=80000, 16K 74.77 53.28 75.11 79.05 63.61 38.20
base=120000, 16K 75.06 53.91 74.85 78.94 63.61 43.60
base=1000000, 16K 74.71 51.95 74.68 79.11 63.30 40.80
@@
Table 3: Short context validation results of scaling RoPE with different base and other extrapolating methods in LLaMA2 (Touvron et al., 2023b). The first box is the result of 7B size and the second box is those of 13B version. LLaMA2 7B or 13B corresponds to the original result as well as that of Dynamic NTK (bloc97, 2023a) since Dynamic NTK does not affect the model within the training context. The final score of SuperGLUE is the average score of all corresponding subtasks.
NQ TriviaQA TruthfulQA ARC-e ARC-c MMLU
LLaMA2 7B 19.06 54.09 35.09 50.26 36.61 46.78
NTK fixed $\alpha=8$ 19.89 53.16 35.09 48.68 34.58 39.36
Linear PI $\lambda=2$ 18.64 52.57 34.80 56.44 35.93 47.27
Linear PI $\lambda=4$ 19.78 52.23 35.09 55.91 36.95 46.80
base=500 20.36 52.43 35.09 53.09 38.64 38.70
base=652 17.12 48.90 34.36 46.74 33.90 33.43
base=1304 17.70 51.27 35.09 49.38 37.63 33.46
base=2608 16.01 51.21 34.06 48.15 34.58 37.07
base=10000 20.14 55.14 35.53 54.32 38.31 47.73
base=40000 17.34 55.27 35.23 55.03 36.95 47.34
base=80000 17.81 54.72 35.23 53.09 37.63 46.87
base=160000 18.98 55.28 35.09 53.26 38.64 46.24
base=400000 20.39 54.91 35.23 52.03 38.64 46.01
base=600000 19.20 54.75 35.67 53.26 37.97 46.39
base=1000000 19.50 54.76 35.82 52.20 37.29 46.29
base=500, 16K 11.58 37.95 34.36 46.21 36.95 28.79
base=10000, 16K 15.29 50.31 34.50 49.91 36.61 42.71
base=40000, 16K 15.32 49.16 34.94 49.38 31.86 45.44
base=80000, 16K 16.51 51.16 34.36 52.20 35.25 45.08
base=120000, 16K 16.15 51.84 34.21 51.85 34.24 43.75
base=1000000, 16K 16.68 53.14 34.80 50.79 34.92 41.83
LLaMA2 13B 24.82 60.69 34.65 59.44 40.00 55.77
NTK fixed $\alpha=8$ 22.77 60.81 34.80 58.38 40.34 52.04
Linear PI $\lambda=2$ 25.96 59.21 34.94 63.49 40.00 57.20
Linear PI $\lambda=4$ 26.07 59.34 35.67 64.20 39.66 56.08
base=500 11.50 57.54 33.77 57.67 38.98 47.98
base=652 15.98 58.67 35.23 58.73 38.31 50.49
base=1304 13.60 59.12 35.23 58.73 41.36 53.59
base=2608 20.00 59.62 34.80 60.67 42.71 55.06
base=10000 26.68 61.57 35.09 62.79 42.03 55.78
base=40000 26.87 60.80 35.53 60.49 42.37 55.26
base=80000 26.76 61.25 35.38 61.20 41.36 55.32
base=160000 27.06 61.22 35.96 60.14 42.03 54.35
base=400000 27.53 60.94 35.82 62.96 42.37 54.01
base=600000 26.81 61.28 35.67 61.20 42.03 53.75
base=1000000 26.51 59.22 35.53 60.85 42.37 52.94
base=500, 16K 11.86 44.74 35.38 49.21 33.22 38.34
base=10000, 16K 19.14 51.22 34.65 58.20 41.02 51.11
base=40000, 16K 23.68 57.32 35.53 59.61 41.69 53.68
base=80000, 16K 24.90 58.74 34.94 61.38 38.98 52.86
base=120000, 16K 25.37 58.16 34.65 61.90 41.02 52.93
base=1000000, 16K 24.35 57.26 35.53 59.44 39.66 51.61
Table 4: Continued table of Table3. The final score of MMLU is the average score of all subtasks.
Next, we validate whether Scaling RoPE, namely the fine-tuning RoPE with different bases on original context length, has a side effect on LLM. We use short context tasks including those listed in the Hugging Face Open LLM Leaderboard(Face, 2023), such as 0-shot Lambada(Paperno et al., 2016), Hellaswag(Zellers et al., 2019), NQ(Kwiatkowski et al., 2019), TriviaQA(Joshi et al., 2017), OBQA(Mihaylov et al., 2018), PIQA(Bisk et al., 2020), TruthfulQA(Lin et al., 2022), Winogrande(Sakaguchi et al., 2020), SuperGLUE(Wang et al., 2019), ARC-easy/challenge(Clark et al., 2018) and 5-shot MMLU(Hendrycks et al., 2021), to assess the performance of RoPE-based LLM after the fine-tuning process. The results obtained are shown in the Table 3 and Table 4. The results of Scaling-RoPE are comparable with that of the pre-trained or directly fine-tuned LLaMA2 (Touvron et al., 2023b). This indicates our approach not only enhances the extrapolation performance but also maintains the inherent knowledge of LLM.
B.3 Long Context Validation
Besides, to validate that our proposed framework can improve the long-range real-world tasks, we also compare the close-ended subtasks included in L-Eval benchmark(An et al., 2023) within 16K context length, including the 16-shot GSM8K(Cobbe et al., 2021), TOEFL-QA(Tseng et al., 2016), TopicRetrieval(Li et al., 2023a), QuALITY(Pang et al., 2022), and Coursera(An et al., 2023). Our method is still capable of these tasks, as shown in Table 5. For RoPE-based LLMs fine-tuned on pre-training context, namely 4K tokens, they still have certain effects on these tasks with 16K context size, which proves one of the claims of our work, that tuning with the original pre-training context can also boost the length extrapolation. This proves that we extrapolate the RoPE-based LLM to a longer context without the disturbance of domain distribution in pre-trained data concerning the change of training context length (Peng et al., 2023). We also report the evaluation results of RoPE-based LLM fine-tuned with longer context, such as 16K tokens, in Table 3, Table 4 and Table 5 and the length extrapolation improvement will be further analyzed in Appendix C.
GSM TOEFL TopicRet QuALITY Coursera
LLaMA2 7B 0.00 34.94 0.00 3.96 0.00
NTK fixed $\alpha=8$ 5.00 34.20 31.33 34.16 21.51
NTK dynamic 9.00 37.17 24.67 30.20 22.09
base=500 4.00 25.65 0.00 18.32 19.19
base=500, log-scaled 3.00 25.65 0.00 18.32 19.77
base=652 6.00 16.73 0.00 13.37 11.63
base=652, log-scaled 6.00 16.73 0.00 16.83 11.63
base=10000 0.00 18.59 0.00 2.97 0.00
base=40000 13.00 34.20 14.67 22.77 11.63
base=80000 11.00 36.43 29.33 23.76 15.70
base=160000 10.00 40.15 38.00 22.28 16.28
base=400000 8.00 41.64 36.67 22.77 15.70
base=600000 12.00 38.66 34.00 25.25 16.86
base=1000000 9.00 31.97 34.00 21.29 16.28
base=2000000 7.00 30.86 31.33 22.77 15.70
base=500, 16K 2.00 25.65 0.00 16.83 20.35
base=10000, 16K 5.00 42.38 0.00 19.80 13.95
base=40000, 16K 11.00 43.49 20.67 28.22 19.19
base=80000, 16K 12.00 43.49 35.33 30.20 21.51
base=120000, 16K 14.00 43.12 40.00 26.24 20.93
base=1000000, 16K 13.00 44.24 41.33 25.74 21.51
LLaMA2 13B 0.00 59.48 0.00 6.44 0.00
NTK fixed $\alpha=8$ 8.00 58.74 36.00 35.64 23.26
NTK dynamic 22.00 59.85 30.00 35.15 28.49
base=500 9.00 37.17 0.00 30.20 25.58
base=500, log-scaled 11.00 37.17 0.00 28.71 25.58
base=652 12.00 37.17 0.00 30.69 25.58
base=652, log-scaled 9.00 37.17 0.00 31.19 25.00
base=10000 0.00 56.13 0.00 7.43 0.00
base=40000 18.00 64.68 20.67 49.01 16.86
base=80000 20.00 63.94 30.00 52.97 24.42
base=160000 23.00 62.83 39.33 51.49 24.42
base=400000 17.00 59.48 34.67 51.49 29.07
base=600000 14.00 63.94 36.00 47.52 29.65
base=1000000 10.00 64.68 28.67 45.05 19.19
base=500, 16K 6.00 20.07 0.00 21.29 26.74
base=10000, 16K 13.00 55.39 0.00 38.61 25.00
base=40000, 16K 25.00 63.94 19.33 49.50 18.60
base=80000, 16K 21.00 63.20 34.67 40.10 15.12
base=120000, 16K 22.00 63.94 40.67 39.11 19.19
base=1000000, 16K 9.00 63.57 33.33 48.02 16.86
Table 5: Long context validation results of scaling RoPE with different base and other extrapolating methods in LLaMA2 (Touvron et al., 2023b). The first box is the result of 7B size and the second box is those of 13B version.
@@ Appendix D Discussion
Besides, we discuss the instructive value of our theory for other extrapolation strategies focused on achieving longer context during the testing phase. These methods are still necessary given two facts. On one hand, the performance of RoPE with a smaller base is still left behind compared with RoPE with much larger bases, such as 1000000, as shown in Figure 1. On the other hand, for RoPE with a base that is not large enough, it still can not extrapolate to a context of 100K or longer as shown in Figure 3. In order to further enhance RoPE’s adaptability to a longer context, whatever the base value is, we discuss the effect of two complementary methods in the inference phase, log-scaled attention (Su, 2023b) and dynamic-scaled RoPE (bloc97, 2023a) on RoPE with different bases.
@@
Besides Log-scaled attention, window method, such as sliding window and its variant, is also a widely accepted strategy for extrapolation, used in inference or evaluation (Press et al., 2022; Sun et al., 2023). Compared with the above strict window-based method, we follow the xPos method proposed in Sun et al. (2023), shown in Equation 17, originally used in the pre-training phase. In this work, we regard this method as a soft sliding window used in the inference phase as use it as a further complement to the log-scaled method. Still, we do little modification besides using the $T_{\text{extra}}$ as the denominator instead of the original denominator $T_{\text{train}}$.
@@
namely Dynamic NTK (bloc97, 2023a) is a widely used extrapolation. Here, we only do two little modifications. One is to change the base 10000 in vanilla RoPE (Su et al., 2021) with the base scaled in the fine-tuning phase, $\beta$. The other is still to replace the $T_{\text{train}}$ in Equation LABEL:equ_ntk with $T_{\text{extra}}$ we derive given Equation 13 in Theorem 3.
$\text{for }\bm{A}_{t,s}\text{, }\theta_{n}={\left(\beta\cdot\alpha_{t}\right)}^{-2n/d}\text{, where }\alpha_{t}=\max\left(1,2^{\left\lceil\log_{2}{\frac{t}{T_{\text{extra}}}}\right\rceil+1}-1\right)\text{.}$ (18)
@@
Figure 11: Perplexity on the validation data from Books3 dataset (Presser, 2020) of LLaMA 7B (Touvron et al., 2023b) based on RoPE with different bases enhanced with log-scaled attention, xPos and dynamic-scaled RoPE and their combination. Here, wrong means calculating $\alpha_{t}$ in Dynamic NTK (bloc97, 2023a) or $p_{t}$ in (Su, 2023b) with training length $T_{\text{train}}$, while correct means using extrapolation upper bound $T_{\text{extra}}$ instead of $T_{\text{train}}$.